# Onsager-Machlup Functional and Large Deviation Principle for Stochastic Hamiltonian Systems

### v1

@@
Abstract.
This paper investigates the application of KAM theory within the probabilistic framework of stochastic Hamiltonian systems. We begin by deriving the Onsager-Machlup functional for the stochastic Hamiltonian system, identifying the most probable transition path for system trajectories. By establishing a large deviation principle for the system, we derive a rate function that quantifies the system’s deviation from the most probable path, particularly in the context of rare events. Furthermore, leveraging classical KAM theory, we demonstrate that invariant tori remain stable, in the most probable sense, despite small random perturbations. Notably, we establish that the exponential decay rate for deviations from these persistent tori coincides precisely with the previously derived rate function, providing a quantitative probabilistic characterization of quasi-periodic motions in stochastic settings.
Key words and phrases:
Stochastic Hamiltonian systems, Onsager-Machlup functional, Large deviation principle, Girsanov transformation, Invariant tori, Euler-Lagrange equation
2020 Mathematics Subject Classification:
Primary 60H10, 82C35, 60F10, 70H08; Secondary 70L05
@@ Abstract.
This work was supported by National Natural Science Foundation of China Grant No. 12071175, 12471183.
1. Introduction
@@
This paper focuses on the persistence of invariant tori in stochastic Hamiltonian systems, particularly examining their stability under random perturbations. We consider the following stochastic Hamiltonian system:
(1.1) $\begin{cases}\,{\rm d}q(t)=\frac{\partial H}{\partial p}(q(t),p(t))\,{\rm d}t+\sigma_{q}(t)\,{\rm d}W_{q}(t),\\ \,{\rm d}p(t)=-\frac{\partial H}{\partial q}(q(t),p(t))\,{\rm d}t+\sigma_{p}(t)\,{\rm d}W_{p}(t),\end{cases}$
@@
where $H(q,p)$ is the Hamiltonian, $\sigma_{q}(t)$ and $\sigma_{p}(t)$ represent the strengths of the random perturbations, and $W_{q}(t)$ and $W_{p}(t)$ are standard Wiener processes. By integrating the Onsager-Machlup functional, large deviation principles and KAM theory, we establish a framework for analyzing the most probable paths and stability of the system under stochastic conditions, thereby revealing the persistence of invariant tori in the most probable context.
@@
Hamiltonian systems are foundational in classical mechanics, with applications across physics, astronomy, mechanical systems, and more. The core framework, dating back to the early 19th century, was introduced by William Rowan Hamilton. Hamiltonian mechanics describes a system’s state via generalized coordinates $q$ and conjugate momenta $p$, while the Hamiltonian function $H(q,p)$ represents the system’s total energy, encompassing both kinetic and potential energy. The evolution of such a system is governed by the Hamiltonian equations:
$\begin{cases}\,{\rm d}q(t)=\frac{\partial H}{\partial p}(q(t),p(t))\,{\rm d}t,\\ \,{\rm d}p(t)=-\frac{\partial H}{\partial q}(q(t),p(t))\,{\rm d}t.\end{cases}$
@@
The original KAM framework was designed primarily for deterministic perturbations, leaving the question of its applicability in stochastic settings unresolved. As a result, validating the extension of the KAM theory under stochastic perturbations and quantifying the stability of invariant tori has emerged as a central problem in stochastic Hamiltonian system research. Recent studies have made significant progress in this area. Wu [44] established a framework for both large and moderate deviations in stochastic Hamiltonian systems, providing a quantitative assessment of the probability of rare events. Talay [42] explored how stochastic Hamiltonian systems asymptotically converge to an invariant measure, with a focus on the exponential nature of this convergence. Lázaro-Camí and Ortega [27] investigated the impact of stochastic noise on classical Hamiltonian dynamics, elucidating how structural preservation can be described in stochastic environments. Zhang [45] examined stochastic flows within Hamiltonian systems and introduced new computational methods based on the Bismut formula. Li [29] proposed an averaging principle framework for completely integrable stochastic Hamiltonian systems, simplifying the analysis of their long-term behavior under small stochastic perturbations. Some of the latest research references are [11, 12, 8] and so on.
The value of this paper lies in its novel analytical framework for understanding the stability of invariant tori in stochastic Hamiltonian systems. By integrating the Onsager-Machlup functional, large deviation theory, and KAM theory, we provide new insights into how these systems behave under stochastic perturbations.
@@
We begin by calculating the Onsager-Machlup functional for stochastic Hamiltonian systems. Due to Hamiltonian system’s symplectic structure, we show that its Onsager-Machlup functional is given by:
$\quad OM(\varphi_{q},\varphi_{p})$
$=\int_{0}^{1}\left\|\sigma_{q}^{-1}(t)\left(\dot{\varphi}_{q}-\frac{\partial H}{\partial\varphi_{p}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}\,{\rm d}t+\int_{0}^{1}\left\|\sigma_{p}^{-1}(t)\left(\dot{\varphi}_{p}+\frac{\partial H}{\partial\varphi_{q}}(\varphi_{q},\varphi_{p})\right)\right\|^{2}\,{\rm d}t.$
Next, we minimize the Onsager-Machlup functional using the Euler-Lagrange equations to determine the most probable continuous path. We demonstrate that this most probable path corresponds to the solution of the deterministic Hamiltonian system without stochastic perturbations. By combining the Onsager-Machlup functional with the Freidlin-Wentzell large deviation theory [28], we establish the large deviation principle for the stochastic Hamiltonian system. Furthermore, we derive the rate function of the system, which quantifies the deviation between rare paths and the most probable path. Finally, leveraging classical KAM theory, we prove that for a nearly integrable Hamiltonian system, the invariant tori remain stable in the most probable sense even under small stochastic perturbations, provided that the system satisfies the non-degeneracy and Diophantine conditions. Importantly, we find that the probability of deviation from these invariant tori can be characterized by the rate function derived earlier.
@@
2. Preliminaries
2.1. Approximate limits in Wiener space
In this section, we recall some fundamental definitions and results concerning approximate limits in Wiener space. Specifically, we focus on the measurable semi-norm, which pertains to the exponentials of random variables in the first and second Wiener chaos (reference [42]).
@@ 2.1. Approximate limits in Wiener space
$\mathbb{H}^{1}$ :=$:=\left\{f:[0,1]\to\mathbb{R}^{n}\in\mathbb{H}^{1}~{}\big{|}~{}f(0)=0,f~{}\text{is absolutely continuous functions}\right.$
$\qquad\left.\text{and}~{}f^{\prime}\in\mathbb{H}\right\}.$
@@
where $\mathcal{P}^{W}$ does not depend on $(h_{1},...,h_{n})$.
@@
We say that a semi-norm $\mathcal{N}$ on $\mathbb{H}^{1}$ is measurable, if there exists a random variable $\tilde{\mathcal{N}}$, satisfying $\tilde{\mathcal{N}}<\infty$ a.s, such that for any approximating sequence of projections $\mathcal{P}_{n}$ on $\mathbb{H}^{1}$, the sequence $\mathcal{N}(\mathcal{P}^{W}_{n})$ converges to $\tilde{\mathcal{N}}$ in probability and $\mathbb{P}(\tilde{\mathcal{N}}\leq\epsilon)>0$ for any $\epsilon>0$. Moreover, if $\mathcal{N}$ is a norm on $\mathbb{H}^{1}$, then we call it a measurable norm.
For proving the measurability of the semi-norm defined in this paper, it is necessary to introduce the following lemma (see [22]).
Lemma 2.3.
Let $\mathcal{N}_{n}$ be a nondecreasing sequence of measurable semi-norms. Suppose that $\tilde{\mathcal{N}}:=\mathbb{P}\text{-}\!\lim\limits_{n\to\infty}\tilde{\mathcal{N}}_{n}$ exists and $\mathbb{P}(\tilde{\mathcal{N}}\leq\epsilon)>0$ for any $\epsilon>0$. In addition, if the limit $\lim\limits_{n\to\infty}\mathcal{N}_{n}$ exists on $\mathbb{H}^{1}$, then $\mathcal{N}:=\lim\limits_{n\to\infty}\mathcal{N}_{n}$ is a measurable semi-norm.
@@
Let $f$ be a function defined on $\Omega$. For $10<\alpha<$ 1, we introduce Hölder norm ($\alpha$-Hölder)
$\|f\|_{\alpha;\Omega}=\|f\|_{0;\Omega}+\left[f\right]_{\alpha;\Omega},$
where $\|f\|_{0;\Omega}$ represents the supremum norm of $f$ on $\Omega$, and $\left[f\right]_{\alpha;\Omega}$ represents the Hölder semi-norm of $f$ on $\Omega$. The specific expression is as follows:
$\|f\|_{0;\Omega}=\sup\limits_{x\in\Omega}|f(x)|,\quad\left[f\right]_{\alpha;\Omega}=\sup\limits_{x,y\in\Omega,x\neq y}\frac{\left|f(x)-f(y)\right|}{|x-y|^{\alpha}}.$
[…]
2.2. Onsager-Machlup functional
@@
Consider a tube surrounding a reference path $\varphi_{t}$ with initial value $\varphi_{0}=x$ and $\varphi_{t}-x$ belongs to $\mathbb{H}^{1}$. Assuming $\epsilon$ is given and small enough, we estimate the probability that the solution process $X_{t}$ is located in that tube as:
@@ 2.3. KAM Theory
We cite the following theorem from [26]:
@@
2.4. Technical Lemmas
In this section, we will introduce several commonly utilized technical lemmas and theorems. Throughout this paper, if not mentioned otherwise, $\mathbb{E}\left(A\big{|}B\right)$ represents the conditional expectation of $A$ under $B$. $C$ represents a positive constant and varies with these different rows.
When we derive the Onsager-Machup functional of SDEs with additive noise, the following lemma is the most basic one, as it ensures that we handle each term separately. Its proof can be found in [24].
Lemma 2.7.
For a fixed integer $N\geq 1$, let $X_{1},...,X_{N}\in\mathbb{R}$ be $N$ random variables defined on $(\Omega,\mathcal{F},\left\{\mathcal{F}_{t}\right\}_{t\geq 0},\mathbb{P})$ and $\left\{D_{\epsilon};\epsilon>0\right\}$ be a family of sets in $\mathcal{F}$. Suppose that for any $c\in\mathbb{R}$ and any $i=1,...,N$,

### v2

# Most Probable KAM Tori in Stochastic Hamiltonian Systems

@@
Abstract
This paper conducts an in-depth investigation into the impact of stochastic perturbations-particularly multiplicative noise-on the integrable structures of Hamiltonian systems, with a central focus on developing a KAM theory for stochastic Hamiltonian dynamics. We begin by deriving the Onsager-Machlup functional for Hamiltonian systems driven by multiplicative noise and identifying the most probable transition paths of system trajectories. This analysis reveals the fundamental differences in how additive versus multiplicative noise influences the integrability of Hamiltonian systems. Building upon this, we establish a large deviation principle for the system and derive a rate function that quantitatively characterizes trajectory deviations, especially in the regime of rare events. The main contribution of this work lies in demonstrating that, under the small noise limit, the quasi-periodic invariant tori of the unperturbed system persist in a probabilistic sense, indicating the stability of KAM structures under stochastic perturbations. Furthermore, we show that the exponential rate of deviation from the invariant tori exactly matches the large deviation rate function, thus providing a quantitative characterization of the structural persistence and fluctuation geometry of quasi-periodic motions in stochastic Hamiltonian systems. These results extend the classical KAM framework to a stochastic setting and offer new insights into the behavior of complex dynamical systems under the influence of noise.
70H08,
60H10,
82C35,
60F10,
70L05,
Stochastic Hamiltonian systems,
Invariant tori,
Onsager-Machlup functional,
Large deviation,
KAM theory,
keywords:
[class=MSC]
keywords:
[…]
, thanks: corresponding author..
Statements and Declarations
  • Ethical approval
    Not applicable.
  • The Data Availability Statement.
    No datasets were generated or analysed during the current study.
  • The Conflict of Interest Statement.
    We have no conflicts of interest to disclose.
  • Funding
@@ Abstract
    The second author (Y. Li) was supported by National Natural Science Foundation of China (Grant No. 12071175, 12471183, 12531009).
1 Introduction
@@
This paper focuses on the persistence of invariant tori in stochastic Hamiltonian systems, particularly examining their stability under random perturbations. We consider the following stochastic Hamiltonian system driven by multiplicative noise:
$\begin{cases}\,{\rm d}\theta(t)=\frac{\partial H}{\partial I}(\theta,I)\,{\rm d}t+\sigma_{\theta}(t,\theta,I)\,{\rm d}W_{\theta}(t),\\ \,{\rm d}I(t)=-\frac{\partial H}{\partial\theta}(\theta,I)\,{\rm d}t+\sigma_{I}(t,\theta,I)\,{\rm d}W_{I}(t),\end{cases}$ (1)
where $H(\theta,I)$ is the Hamiltonian, $\theta$ represents the angle variables, and $I$ represents the action variables. $\sigma_{\theta}(t,\theta,I)$ and $\sigma_{I}(t,\theta,I)$ represent the strengths of the random perturbations, and $W_{\theta}(t)$ and $W_{I}(t)$ are standard Wiener processes. The corresponding symplectic form of this system is given by
$\mathrm{d}X(t)=J\nabla H(X(t))\,\mathrm{d}t+\sigma(t,X(t))\,\mathrm{d}W(t),$ (2)
where $J$ is the standard symplectic matrix defined by
$J:=\begin{pmatrix}0&I_{n}\\ -I_{n}&0\end{pmatrix}\in\mathbb{R}^{2n\times 2n},$
and $\nabla H(X)$ denotes the gradient of the Hamiltonian with respect to $X=(\theta,I)$, given by
$\nabla H(X)=\begin{pmatrix}\frac{\partial H}{\partial\theta}\\ \frac{\partial H}{\partial I}\end{pmatrix}$,
and the diffusion matrix $\sigma(t,X)$ is defined as the block-diagonal matrix
$\sigma(t,X):=\begin{pmatrix}\sigma_{\theta}(t,\theta,I)&0\\ 0&\sigma_{I}(t,\theta,I)\end{pmatrix}\in\mathbb{R}^{2n\times 2n}.$
@@
By integrating the Onsager-Machlup functional, large deviation principles, and KAM theory, we establish a framework for analyzing the most probable paths and stability of the system under stochastic conditions, thereby revealing the persistence of invariant tori in the most probable context.
@@
Hamiltonian systems are foundational in classical mechanics, with applications across physics, astronomy, mechanical systems, and more. The core framework, dating back to the early 19th century, was introduced by William Rowan Hamilton. Hamiltonian mechanics describes a system’s state via generalized coordinates $\theta$ and conjugate momenta $I$, while the Hamiltonian function $H(\theta,I)$ represents the system’s total energy, encompassing both kinetic and potential energy. The evolution of such a system is governed by the Hamiltonian equations
$\mathrm{d}X(t)=J\nabla H(X(t))\,\mathrm{d}t.$
@@
The original KAM framework was designed primarily for deterministic perturbations, leaving the question of its applicability in stochastic settings unresolved. As a result, validating the extension of the KAM theory under stochastic perturbations and quantifying the stability of invariant tori has emerged as a central problem in stochastic Hamiltonian system research. Recent studies have made significant progress in this area. Wu [46] established a framework for both large and moderate deviations in stochastic Hamiltonian systems, providing a quantitative assessment of the probability of rare events. Talay [44] explored how stochastic Hamiltonian systems asymptotically converge to an invariant measure, with a focus on the exponential nature of this convergence. Lázaro-Camí and Ortega [27] investigated the impact of stochastic noise on classical Hamiltonian dynamics, elucidating how structural preservation can be described in stochastic environments. Li [30] proposed an averaging principle framework for completely integrable stochastic Hamiltonian systems, simplifying the analysis of their long-term behavior under small stochastic perturbations. Zhang [47] examined stochastic flows within Hamiltonian systems and introduced new computational methods based on the Bismut formula. Some of the latest research references are [29, 11, 8] and so on.
The value of this paper lies in the development of a novel analytical framework for investigating the stability of invariant tori in stochastic Hamiltonian systems. By integrating the Onsager-Machlup functional, large deviation theory, and KAM theory, we provide new insights into the behavior of such systems under stochastic perturbations, particularly in the presence of multiplicative noise. The main contributions of this work are summarized as follows:
  1. 1.
    We establish, for the first time, a rigorous formulation of the Onsager-Machlup functional for stochastic Hamiltonian systems with multiplicative noise, thereby extending previous results which were restricted to additive noise. Our formulation accommodates general nonlinear diffusion coefficients that depend on the state variables and captures the non-Gaussian nature of path distributions. To overcome technical challenges arising from non-adapted path perturbations, we further develop a conditional expectation comparison principle and a small ball probability estimate tailored for the multiplicative noise setting.
  2. 2.
    Using the Ansatz method, we rigorously prove that the most probable path satisfies a modified deterministic Hamiltonian equation. On this basis, we derive an explicit large deviation principle and the corresponding rate functional via the Freidlin–Wentzell framework, providing a quantitative characterization of the probability that rare trajectories deviate from the invariant tori.
  3. 3.
    We extend the classical KAM theory to the stochastic setting, and prove that under the most probable path interpretation, invariant tori of nearly integrable Hamiltonian systems remain stable under small stochastic perturbations, provided the non-degeneracy and Diophantine conditions are satisfied. This result offers a new probabilistic perspective on the long-term stability of Hamiltonian systems.
@@
A major difficulty in this work arises from the more challenging multiplicative noise setting, in which the diffusion coefficient $\sigma(t,X_{t})$ depends nonlinearly on the current state $X_{t}$. This leads to non-Gaussian path distributions and introduces significant analytical challenges. Specifically, the classical Cameron–Martin transformation no longer preserves the law of the process, making small ball estimates and conditional expectations difficult to control using standard Gaussian techniques. Moreover, path perturbations no longer satisfy linear superposition, rendering traditional tools such as orthogonal expansions and characteristic function methods inapplicable. The lack of explicit density formulas for non-Gaussian paths further complicates the analysis of conditional expectations and path-dependent functionals.
Notably, foundational results in Onsager-Machlup theory—such as Theorem 2.12 and Theorem 2.14—fail to hold under such general settings. To address these challenges, we first establish a small ball probability estimate (Lemma 2.19), which serves as a key technical tool. Building on this result, we introduce a conditional expectation comparison principle, which extends Theorems 2.12 and 2.14, leading to Lemmas 2.20 and 2.21 that are tailored for the multiplicative noise framework.
On this foundation, we rigorously characterize the Onsager-Machlup functional for the multiplicative noise system using advanced non-Gaussian techniques, including the Radon-Nikodym derivative structure under Girsanov’s transformation, non-adapted Skorohod integration, small ball probability estimates, and conditional expectation bounds. We show that the Onsager-Machlup functional is given by
$OM(\varphi,\dot{\varphi})=\int_{0}^{1}\left\|\sigma^{-1}(t,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right\|^{2}\,\mathrm{d}s-\int_{0}^{1}\mathrm{Tr}\left(\left(D_{y}\sigma(s,\varphi_{s})\right)\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right)\,\mathrm{d}s.$
Next, we apply the Ansatz method and the variational approach to minimize the Onsager-Machlup functional and determine the most probable continuous path of the system. We prove that this most probable path corresponds to the solution of the modified deterministic Hamiltonian equation
$\dot{\varphi}(t)=J\nabla H(\varphi_{t})-\frac{1}{2}\sigma(t,\varphi_{t})\sum_{i}\left(D_{y}\sigma(t,\varphi_{t})\right)_{\cdot ii}.$
By combining the Onsager-Machlup functional with the Freidlin-Wentzell large deviation theory [28], we establish a large deviation principle for the stochastic Hamiltonian system. Furthermore, we derive the associated rate function
$S(\varphi)=\frac{1}{2}\int_{0}^{1}\left\|\sigma^{-1}(s,\varphi_{s})\left(J\nabla H(\varphi_{s})-\dot{\varphi}_{s}\right)\right\|^{2}\,\mathrm{d}s,$
which quantifies the deviation of rare paths from the most probable trajectory. We observe that this rate function is closely related to the Onsager-Machlup functional, with the trace-related term in the latter corresponding to higher-order contributions in the small noise limit and hence vanishing under the large deviation scaling.
Finally, by employing the classical KAM theory, we prove that for a nearly integrable Hamiltonian system, the invariant tori remain stable in the most probable sense under small stochastic perturbations, provided the non-degeneracy and Diophantine conditions are satisfied. Notably, we find that the probability of deviation from these invariant tori is precisely characterized by the rate function derived above.
@@
2 Preliminaries
2.1 Basic space and norm
In this section, we review some fundamental definitions and results concerning asymptotic limits in the Wiener space (see reference [44]), as well as several basic definitions of functional spaces and norms (see reference [20, 22]).
@@ 2.1 Basic space and norm
$\mathbb{H}^{1}$ $=\left\{f:[0,1]\to\mathbb{R}^{n}\in\mathbb{H}^{1}~\big{|}~f(0)=0,f~\text{is absolutely continuous functions and}~f^{\prime}\in\mathbb{H}\right\}.$
@@
where $\mathcal{P}^{W}$ is independent of the choice of the orthonormal basis $(h_{1},\ldots,h_{n})$.
@@
We say that a semi-norm $\mathcal{N}$ on $\mathbb{H}^{1}$ is measurable, if there exists a random variable $\tilde{\mathcal{N}}$, satisfying $\tilde{\mathcal{N}}$$\infty$ a.s, such that for any approximating sequence of projections $\mathcal{P}_{n}$ on $\mathbb{H}^{1}$, the sequence $\mathcal{N}(\mathcal{P}^{W}_{n})$ converges to $\tilde{\mathcal{N}}$ in probability and $\mathbb{P}(\tilde{\mathcal{N}}\leq\epsilon)>0$ for any $\epsilon>0$. Moreover, if $\mathcal{N}$ is a norm on $\mathbb{H}^{1}$, then we call it a measurable norm.
@@
We define the space $C^{2}_{b}([0,1],\mathbb{R}^{n\times n})$ as the set of matrix-valued functions $g(t)$ that satisfy the following conditions:
  1. 1.
    For every $t\in[0,1]$, $g(t)$ is a real symmetric matrix;
  2. 2.
    $g(t)$ is twice continuously differentiable;
  3. 3.
    $g(t)$ is bounded, meaning there exists a constant $M>0$ such that $\|g(t)\|_{2}\leq M$ for all $t\in[0,1]$, where $\|g(t)\|_{2}$ denotes the spectral norm of the matrix $g(t)$.
Definition 2.4.
We define the space $C^{2}_{b}([0,1]\times\mathbb{R}^{d},\mathbb{R}^{n\times n})$ as the set of matrix-valued functions $g(t,x)$ that satisfy the following conditions:
  1. 1.
    For every $(t,x)\in[0,1]\times\mathbb{R}^{d}$, $g(t,x)$ is a real symmetric matrix;
  2. 2.
    $g(t,x)$ is twice continuously differentiable with respect to all variables $(t,x)$;
  3. 3.
    $g(t,x)$ is bounded, meaning there exists a constant $M>0$ such that $\|g(t,x)\|_{2}\leq M$ for all $(t,x)\in[0,1]\times\mathbb{R}^{d}$, where $\|g(t,x)\|_{2}$ denotes the spectral norm of the matrix $g(t,x)$.
Definition 2.5.
For a matrix $A=[a_{ij}]\in\mathbb{R}^{n\times n}$, its spectral norm is defined as:
$\|A\|_{2}=\max_{\|x\|_{2}=1}\|Ax\|_{2}.$
Alternatively, it can be expressed as:
$\|A\|_{2}=\sigma_{\text{max}}(A),$
where $\sigma_{\text{max}}(A)$ denotes the largest singular value of the matrix $A$. The singular values of $A$ are the square roots of the eigenvalues of the matrix $A^{T}A$.
Definition 2.6.
Let $A=[a_{ij}]\in\mathbb{R}^{n\times n}$. Its Hilbert–Schmidt norm (also known as the Frobenius norm) is given by
$\|A\|_{\mathrm{HS}}\;=\;\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{n}\bigl{\lvert}a_{ij}\bigr{\rvert}^{2}}\;=\;\sqrt{\mathrm{trace}\bigl{(}A^{T}A\bigr{)}}.$
Equivalently, if $\{\sigma_{k}(A)\}$ are the singular values of $A$, then
$\|A\|_{\mathrm{HS}}\;=\;\sqrt{\sum_{k=1}^{n}\sigma_{k}(A)^{2}}.$
Thus the Hilbert–Schmidt norm captures the sum of the squares of all entries (or, equivalently, the square root of the sum of the squares of all singular values) of the matrix.
Definition 2.7.
Let $f$ be a function defined on a compact domain $\Omega\subset\mathbb{R}^{d}$. The supremum norm (also known as the maximum norm) of $f$ on $\Omega$ is defined by
$\|f\|_{0;\Omega}:=\sup_{x\in\Omega}|f(x)|,$
where $|\cdot|$ denotes the Euclidean norm in $\mathbb{R}^{d}$. This norm quantifies the largest absolute value attained by $f$ over the domain $\Omega$.
2.2 Onsager-Machlup functional
@@
Consider a tube surrounding a reference path $\varphi_{t}$ with initial value $\varphi_{0}=x$ and $\varphi_{t}-x$ belongs to $\mathbb{H}^{1}$. Assuming $\epsilon$ is given and small enough, we estimate the probability that the solution process $X_{t}$ is located in that tube as
@@ 2.3 KAM Theory
We will use the following theorem, which is stated as Theorem 3.1 in [26].
@@
2.4 Technical theorems
In this section, we will introduce several commonly utilized technical theorems. Throughout this paper, unless otherwise specified, the norm $\|\cdot\|$ refers to the supremum norm $\|\cdot\|_{0;\Omega}$. We use $C$ to denote a generic constant, which may vary from line to line. The symbol $n$ denotes the spatial dimension, which is assumed to be an arbitrary but finite positive integer. The notation $\mathbb{E}(A\mid B)$ stands for the conditional expectation of the random variable $A$ given the event or $\sigma$-algebra $B$.
When we derive the Onsage-Machup functional of stochastic differential equations with additive noise, the following theorem is the most basic one, as it ensures that we handle each term separately. For a complete proof, see [23], pages 536-537.
Theorem 2.10.
For a fixed integer $N\geq 1$, let $X_{1},...,X_{N}\in\mathbb{R}$ be $N$ random variables defined on $(\Omega,\mathcal{F},\left\{\mathcal{F}_{t}\right\}_{t\geq 0},\mathbb{P})$ and $\left\{D_{\epsilon};\epsilon>0\right\}$ be a family of sets in $\mathcal{F}$. Suppose that for any $c\in\mathbb{R}$ and any i=1,,Ni=1,...,N, we have